# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 4 Existence for evolution problem
(4.53) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ \hskip 85.35826pt-\int_{0}^{t}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ for any }t\in[0,T).\end{array}$
@@
Remember that $\frac{\eta_{1}^{i}}{\alpha_{1}}\vee\frac{\eta_{2}^{i}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}^{i}+\alpha_{2}\rho_{2}^{i}).$ This implies that $(\rho^{i}-\rho^{i-1})\cdot\eta^{i}\geq\beta_{\alpha}[\rho^{i}]-\beta_{\alpha}[\rho^{i-1}],$ and then
(4.55) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{i}]-\rho^{i}\cdot\tilde{g})-\int_{\Omega}(\beta_{\alpha}[\rho^{i-1}]-\rho^{i-1}\cdot\tilde{g})+\tau\>\int_{\Omega}F_{\sigma}[\nabla(\eta^{i}-\tilde{g})]\leq\tau\>[f^{i},\eta^{i}-\tilde{g}]_{\Omega}\\ \\ -\tau\>\int_{\Omega}[\rho^{i-1}\>V^{i}\cdot\nabla(\eta^{i}-\tilde{g})]+\tau\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\tau\>[\pi,\eta^{i}]_{\Gamma_{N}}.\end{array}$
@@
(4.56) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T),\end{array}$
@@
(4.57) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\int_{0}^{t_{l}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t_{l}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\int_{0}^{t_{l}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t_{l}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T).\end{array}$
@@
(4.58) $\begin{array}[]{c}\int_{\Omega}(\alpha_{1}\>\rho_{1}^{\tau}(t)+\alpha_{2}\>\rho_{2}^{\tau}(t)-M)^{+2}+\int_{0}^{T}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\\ \\ \leq\frac{B}{C}\left\{1+\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\>\exp\left(\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\right)\right\},\end{array}$
@@
$B=\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}(\rho_{01}+\rho_{02}-M)^{+2}$
@@
(4.59) $\begin{array}[]{c}\int_{\Omega}(\beta_{\alpha}[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]\\ \\ +\underbrace{\int_{0}^{t}\!\!\int_{\Omega}|[\rho^{\tau}(.-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]|}_{J_{1}^{\tau}}+\underbrace{\int_{0}^{t}|[\pi,\eta^{\tau}]_{\Gamma_{N}}+[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}|}_{J_{2}^{\tau}}.\end{array}$
@@
$\displaystyle J_{1}^{\tau}$$\leq$ $\displaystyle C_{\epsilon}\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}(.-\tau)+\alpha_{2}\rho_{2}(.-\tau)-M)^{+2}+\epsilon\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]$
@@
$\displaystyle+\tau\>\int_{\Omega}(\alpha_{1}\rho_{01}+\alpha_{2}\rho_{02}-M)^{+2}.$
@@
$\beta_{\alpha}[\rho^{\tau}]-\rho^{\tau}\cdot\tilde{g}\geq\frac{C}{2}(\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}_{1}\vee\tilde{g}_{2}|^{2}\geq\frac{C}{2}(\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}|^{2},$
@@
(4.60) $\begin{array}[]{c}\frac{C}{2}\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}+(1-2\epsilon)\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\frac{1}{2C}\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+T\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]\\ \\ +\int_{\Omega}(\beta_{\alpha}[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\tau\>\int_{\Omega}(\alpha_{1}\rho_{01}+\alpha_{2}\rho_{02}-M)^{+2}+C_{\epsilon}\>\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}-M)^{+2}\\ \\ +\int_{0}^{T}\!\!\left(\|\pi\|_{[H^{-1/2}_{\Gamma_{N}}]^{2}}^{2}+\|f_{0}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}+\|\overline{f}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}\right)+C_{\epsilon}\>M^{2}\>T\>\|V^{\tau}\|_{\infty}^{2}|\Omega|,\quad\hbox{for any }t\in[0,T).\end{array}$
@@
(4.61) $\begin{array}[]{c}\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla\eta^{\tau}-\tilde{g}]\\ \\ \leq\frac{B}{C}+\frac{\|V^{\tau}\|_{\infty}^{2}}{C}\>\int_{0}^{t}\!\!\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}.\end{array}$
@@
(4.62) $\int_{\Omega}(\alpha_{1}\rho^{\tau}_{1}(t)+\alpha_{2}\rho^{\tau}_{2}(t)-M)^{+2}\leq\frac{B}{C}\>\exp\left(T\|V^{\tau}\|_{\infty}^{2}/C\right),$
@@
with $(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}$ on $\Sigma_{N_{k}},$ and by (4.58), $\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho_{k}^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau}$ is bounded in $L^{2}(Q),$ we see that $\partial_{t}\tilde{\rho}_{k}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)\right).$ Combining this with (4.47) we deduce that $\tilde{\rho}^{\tau}$ and $\rho^{\tau}$ have the same $[L^{2}(Q)]-$weak limit. Thus (4.50). Let us prove now that the couple $(\rho,\eta)$ satisfies the state equations (2.14). Thanks again to (4.58), $\eta_{1}^{\tau}\vee\eta_{2}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{1}(\Omega)\right),$ and there exists $\overline{\eta}\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right)$ such that, by taking a sub-sequence if necessary,
(4.63) $\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\to\overline{\eta},\quad\hbox{ in }L^{2}\left(0,T;H^{1}(\Omega)\right)-\hbox{weak},$
which satisfies moreover
(4.64) $0\leq\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\leq\overline{\eta},\quad\hbox{ a.e. in }Q.$
So, using Weak Aubin’s type Lemma (cf. [24, 5]), we get
(4.65) $\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}^{\tau}\>\eta_{k}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}\>\eta_{k}\>\varphi,\quad\forall\>\varphi\in\mathcal{D}(Q)\hbox{ and }k=1,2,$
@@
(4.66) $\int_{0}^{T}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau})\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\overline{\eta}\>\varphi\quad\forall\>\varphi\in\mathcal{D}(Q).$
Combining this with the fact that $\rho_{1}^{\tau}\>\eta_{1}^{\tau}+\rho_{2}^{\tau}\>\eta_{2}^{\tau}=(\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau})\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}},$ we obtain
(4.67) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\overline{\eta},\quad\hbox{ a.e. in }Q,$
and, by (4.64), we deduce that
(4.68) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}},\quad\hbox{ a.e. in }Q.$
Moreover, since $\frac{\eta_{1}^{\tau}}{\alpha_{1}}\vee\frac{\eta_{2}^{\tau}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}^{\tau}+\alpha_{2}\rho_{2}^{\tau}),$ a.e. in $Q,$ by using again weak Aubin’s type Lemma and (4.66), we see that
$\overline{\eta}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),\quad\hbox{ a.e. in }Q.$
To finish the proof, let us justify that $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ a.e. in $Q.$ If $\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}\neq 0,$ (4.67) and (4.68) imply clearly that $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}=\overline{\eta}\in\partial\beta(.,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ a.e. in $Q.$ If $\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}=0,$ the result follows from the fact that $\partial\beta(.,0)\ni 0\leq\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}2\leq\overline{\eta}\in\partial\beta(.,0).$

### v2

@@ 4 Existence for evolution problem
(4.68) $\begin{array}[]{c}\int_{\Omega}(\beta[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ \hskip 85.35826pt-\int_{0}^{t}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ for any }t\in[0,T).\end{array}$
@@
Remember that $\ \eta_{1}^{i}\vee\eta_{2}^{i}\in\partial\beta(.,\rho_{1}^{i}+\rho_{2}^{i}).$ This implies that $(\rho^{i}-\rho^{i-1})\cdot\eta^{i}\geq\beta[\rho^{i}]-\beta[\rho^{i-1}],$ and then
(4.70) $\begin{array}[]{c}\int_{\Omega}(\beta[\rho^{i}]-\rho^{i}\cdot\tilde{g})-\int_{\Omega}(\beta[\rho^{i-1}]-\rho^{i-1}\cdot\tilde{g})+\tau\>\int_{\Omega}F_{\sigma}[\nabla(\eta^{i}-\tilde{g})]\leq\tau\>[f^{i},\eta^{i}-\tilde{g}]_{\Omega}\\ \\ -\tau\>\int_{\Omega}[\rho^{i-1}\>V^{i}\cdot\nabla(\eta^{i}-\tilde{g})]+\tau\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\tau\>[\pi,\eta^{i}]_{\Gamma_{N}}.\end{array}$
@@
(4.71) $\begin{array}[]{c}\int_{\Omega}(\beta[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]+\sum_{i=1}^{l}\int_{t_{i-1}}^{t_{i}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T),\end{array}$
@@
(4.72) $\begin{array}[]{c}\int_{\Omega}(\beta[\rho^{\tau}(t_{l})]-\rho^{\tau}(t_{l})\cdot\tilde{g})+\int_{0}^{t_{l}}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\int_{0}^{t_{l}}[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}\\ \\ -\int_{0}^{t_{l}}\!\!\int_{\Omega}[\rho^{\tau}(t-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]+T\>\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{0}^{t_{l}}[\pi,\eta^{\tau}]_{\Gamma_{N}},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T).\end{array}$
@@
(4.73) $\begin{array}[]{c}\int_{\Omega}(\rho_{1}^{\tau}(t)+\rho_{2}^{\tau}(t)-M)^{+2}+\int_{0}^{T}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\\ \\ \leq\frac{B}{C}\left\{1+\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\>\exp\left(\frac{T\|V^{\tau}\|_{\infty}^{2}}{C}\right)\right\},\end{array}$
@@
$B=\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]+\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}(\rho_{01}+\rho_{02}-M)^{+2}$
@@
(4.74) $\begin{array}[]{c}\int_{\Omega}(\beta[\rho^{\tau}(t)]-\rho^{\tau}(t)\cdot\tilde{g})+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+T\>\int_{\Omega}F_{\sigma}[\sigma\>\nabla\tilde{g}]\\ \\ +\underbrace{\int_{0}^{t}\!\!\int_{\Omega}|[\rho^{\tau}(.-\tau)V^{\tau}\cdot\nabla(\eta^{\tau}-\tilde{g})]|}_{J_{1}^{\tau}}+\underbrace{\int_{0}^{t}|[\pi,\eta^{\tau}]_{\Gamma_{N}}+[f^{\tau},\eta^{\tau}-\tilde{g}]_{\Omega}|}_{J_{2}^{\tau}}.\end{array}$
@@
$\displaystyle J_{1}^{\tau}$$\leq$ $\displaystyle C_{\epsilon}\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\rho_{1}(.-\tau)+\rho_{2}(.-\tau)-M)^{+2}+\epsilon\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]$
@@
$\displaystyle+\tau\>\int_{\Omega}(\rho_{01}+\rho_{02}-M)^{+2}.$
@@
$\beta[\rho^{\tau}]-\rho^{\tau}\cdot\tilde{g}\geq\frac{C}{2}(\rho^{\tau}_{1}+\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}_{1}\vee\tilde{g}_{2}|^{2}\geq\frac{C}{2}(\rho^{\tau}_{1}+\rho^{\tau}_{2}-M)^{+2}-\frac{1}{2C}|\tilde{g}|^{2},$
@@
(4.75) $\begin{array}[]{c}\frac{C}{2}\int_{\Omega}(\rho^{\tau}_{1}(t)+\rho^{\tau}_{2}(t)-M)^{+2}+(1-2\epsilon)\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla(\eta^{\tau}-\tilde{g})]\leq\frac{1}{2C}\|\tilde{g}\|_{[L^{2}(\Omega)]^{2}}^{2}+T\int_{\Omega}F_{\sigma}[\nabla\tilde{g}]\\ \\ +\int_{\Omega}(\beta[\rho_{0}]-\rho_{0}\cdot\tilde{g})+\tau\>\int_{\Omega}(\rho_{01}+\rho_{02}-M)^{+2}+C_{\epsilon}\>\|V^{\tau}\|_{\infty}^{2}\int_{0}^{t}\!\!\int_{\Omega}(\rho_{1}+\rho_{2}-M)^{+2}\\ \\ +\int_{0}^{T}\!\!\left(\|\pi\|_{[H^{-1/2}_{\Gamma_{N}}]^{2}}^{2}+\|f_{0}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}+\|\overline{f}^{\tau}\|_{[L^{2}(\Omega)]^{2}}^{2}\right)+C_{\epsilon}\>M^{2}\>T\>\|V^{\tau}\|_{\infty}^{2}|\Omega|,\quad\hbox{for any }t\in[0,T).\end{array}$
@@
(4.76) $\begin{array}[]{c}\int_{\Omega}(\rho^{\tau}_{1}(t)+\rho^{\tau}_{2}(t)-M)^{+2}+\int_{0}^{t}\!\!\int_{\Omega}F_{\sigma}[\nabla\eta^{\tau}-\tilde{g}]\\ \\ \leq\frac{B}{C}+\frac{\|V^{\tau}\|_{\infty}^{2}}{C}\>\int_{0}^{t}\!\!\int_{\Omega}(\rho^{\tau}_{1}(t)+\rho^{\tau}_{2}(t)-M)^{+2}.\end{array}$
@@
(4.77) $\int_{\Omega}(\rho^{\tau}_{1}(t)+\rho^{\tau}_{2}(t)-M)^{+2}\leq\frac{B}{C}\>\exp\left(T\|V^{\tau}\|_{\infty}^{2}/C\right),$
@@
with $(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}$ on $\Sigma_{N_{k}},$ and by (4.73), $\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho_{k}^{\tau}(.-\tau)V_{k}^{\tau}+\overline{f}_{k}^{\tau}$ is bounded in $L^{2}(Q),$ we see that $\partial_{t}\tilde{\rho}_{k}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)\right).$ Combining this with (4.61) we deduce that $\tilde{\rho}^{\tau}$ and $\rho^{\tau}$ have the same $[L^{2}(Q)]-$weak limit. Thus (4.64) and (4.65). Let us prove now that the couple $(\rho,\eta)$ satisfies the state equations (2.17). Thanks again to (4.73), we see that $\eta_{1}^{\tau}\vee\eta_{2}^{\tau}$ is bounded in $L^{2}\left(0,T;H^{1}(\Omega)\right),$ and there exists $\overline{\eta}\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right)$ such that
(4.78) $0\leq{\eta_{1}}\vee\eta_{2}\leq\overline{\eta},\quad\hbox{ a.e. in }Q,$
and, by taking a sub-sequence if necessary,
(4.79) $\eta_{1}^{\tau}\vee\eta_{2}^{\tau}\to\overline{\eta},\quad\hbox{ in }L^{2}\left(0,T;H^{1}(\Omega)\right)-\hbox{weak},$
Now, using weak compensated compactness results (cf. [24, 5]), we get
(4.80) $\int_{0}^{T}\!\!\int_{\Omega}\tilde{\rho}_{k}^{\tau}\>\eta_{k}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}\>\eta_{k}\>\varphi,\hbox{ and }k=1,2,$
@@
(4.81) $\int_{0}^{T}\!\!\int_{\Omega}(\tilde{\rho}_{1}^{\tau}+\tilde{\rho}_{2}^{\tau})\>\eta_{1}^{\tau}\vee\eta_{2}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}(\rho_{1}+\rho_{2})\>\overline{\eta}\>\varphi,$
for any $\varphi\in\mathcal{D}(Q).$ Combining this with (4.61) and the fact that $\partial_{t}\tilde{\rho}^{\tau}$ and $\eta^{\tau}$ are bounded in $L^{2}\left(0,T;H^{-1}_{\Gamma_{D_{k}}}(\Omega)\right)$ and $L^{2}\left(0,T;H^{1}(\Omega)\right),$ respectively, we deduce that
(4.82) $\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}^{\tau}\>\eta_{k}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}\rho_{k}\>\eta_{k}\>\varphi,\hbox{ and }k=1,2,$
and
(4.83) $\int_{0}^{T}\!\!\int_{\Omega}(\rho_{1}^{\tau}+\rho_{2}^{\tau})\>\eta_{1}^{\tau}\vee\eta_{2}^{\tau}\>\varphi\quad\xrightarrow[\tau\to 0]{}\quad\int_{0}^{T}\!\!\int_{\Omega}(\rho_{1}+\rho_{2})\>\overline{\eta}\>\varphi,$
for any $\varphi\in\mathcal{D}(Q).$ Moreover, since $\eta_{1}^{\tau}\vee\eta_{2}^{\tau}\in\partial\beta(.,\rho_{1}^{\tau}+\rho_{2}^{\tau}),$ a.e. in $Q,$ by using Weak Aubin’s type Lemma (cf. Proposition 1.4 in [5]), we see that
$\overline{\eta}\in\partial\beta(.,\rho_{1}+\rho_{2}),\quad\hbox{ a.e. in }Q.$
To finish the proof, let us justify that ${\eta_{1}}\vee\eta_{2}\in\partial\beta(.,\rho_{1}+\rho_{2}),$ a.e. in $Q.$ First, one sees that since $(\rho_{1}^{\tau}+\rho_{2}^{\tau})\>\eta_{1}^{\tau}\vee\eta_{2}^{\tau}=\rho_{1}^{\tau}\>\eta_{1}^{\tau}+\rho_{2}^{\tau}\>\eta_{2}^{\tau},$ (4.82) and (4.83) imply
(4.84) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\rho_{1}+\rho_{2})\>\overline{\eta},\quad\hbox{ a.e. in }Q,$
and, using (4.78), we deduce that
(4.85) $\rho_{1}\>\eta_{1}+\rho_{2}\>\eta_{2}=(\rho_{1}+\rho_{2})\>{\eta_{1}}\vee\eta_{2},\quad\hbox{ a.e. in }Q.$
So, if $\rho_{1}+\rho_{2}\neq 0,$ (4.84) and (4.85) imply clearly that ${\eta_{1}}\vee\eta_{2}=\overline{\eta}\in\partial\beta(.,\rho_{1}+\rho_{2}),$ a.e. in $Q.$ And, if $\rho_{1}+\rho_{2}=0,$ the result follows from the fact that $\partial\beta(.,0)\ni 0\leq{\eta_{1}}\vee\eta_{2}\leq\overline{\eta}\in\partial\beta(.,0).$
@@
Remark 7
Thanks to compensated compactness results of Proposition 1.4 in [5], if $\beta^{-1}(x,.)$ is single valued then we have $\rho_{1}^{\tau}+\rho_{2}^{\tau}\to\rho_{1}+\rho_{2}$ in $L^{2}(Q).$ However, it is not clear if this remains to be true for each density $\rho_{k}.$